NOAA Office of Satellite and Product Operations

NOAA 15 HIRS Subsystem Summary

Component Description Status
CHANNEL 1 Spectral Channel #1, 14.95 micro meters Green
CHANNEL 10 Spectral Channel #10, 12.47 micro meters Green
CHANNEL 11 Spectral Channel #11, 7.33 micro meters Green
CHANNEL 12 Spectral Channel #12, 6.52 micro meters Green
CHANNEL 13 Spectral Channel #13, 4.57 micro meters Green
CHANNEL 14 Spectral Channel #14, 4.52 micro meters Yellow
CHANNEL 15 Spectral Channel #15, 4.47 micro meters Green
CHANNEL 16 Spectral Channel #16, 4.45 micro meters Yellow
CHANNEL 17 Spectral Channel #17, 4.13 micro meters Yellow
CHANNEL 18 Spectral Channel #18, 4.00 micro meters Green
CHANNEL 19 Spectral Channel #19, 3.76 micro meters Green
CHANNEL 2 Spectral Channel #2, 14.71 micro meters Green
CHANNEL 20 Spectral Channel #20, 0.69 micro meters Green
CHANNEL 3 Spectral Channel #3, 14.49 micro meters Green
CHANNEL 4 Spectral Channel #4, 14.22 micro meters Green
CHANNEL 5 Spectral Channel #5, 13.97 micro meters Green
CHANNEL 6 Spectral Channel #6, 13.64 micro meters Green
CHANNEL 7 Spectral Channel #7, 13.35 micro meters Green
CHANNEL 8 Spectral Channel #8, 11.11 micro meters Green
CHANNEL 9 Spectral Channel #9, 9.71 micro meters Green
ELECTRONICS Instrument Electronics Green
FILTER WHEEL Instrument Filter Wheel Motor Red
SCAN MOTOR Instrument Scan Motor Green

Subsystem History:

Date Time Component Description
01/10/2012 16:15:09 FILTER WHEEL Motor Power Toggling Macro (once per orbit) restarted.
10/03/2011 16:06:04 FILTER WHEEL Disabled Filter Wheel Power Toggle Macros
04/22/2010 19:51:00   Disabling the HIRS Filter Motor Housing Heater macros during the warm summer season.
04/20/2010 14:16:00   Load MACRO20 and MACRO21 then start MACRO20 - Filter Motor Housing Heater Power Toggling Macros
06/22/2009 16:42:30 FILTER WHEEL HIRS Filter Wheel Housing Heater On
06/22/2009 16:33:53 FILTER WHEEL Toggled the Filter Wheel Motor On/Off 30 times from 16:34-16:42 in attempts to move the stalled motor. Motor left On at end of sequence.
06/19/2009 17:54:58 FILTER WHEEL HIRS Filter Wheel Housing Heater Off
06/19/2009 17:54:24 FILTER WHEEL HIRS Filter Wheel Motor Power Off
06/19/2009 17:47:26 FILTER WHEEL Toggled the Filter Wheel Motor On/Off 18 times from 17:47-17:53 in attempts to move the stalled motor.
06/19/2009 14:27:12 FILTER WHEEL Toggled the Filter Wheel Motor On/Off 7 times from 14:27-14:33 in attempts to move the stalled motor.
06/16/2009 17:19:33 FILTER WHEEL HIRS Filter Wheel to High Power
06/05/2009 19:46:40 FILTER WHEEL HIRS Filter Motor Power to Normal
06/05/2009 16:43:35 FILTER WHEEL HIRS Filter Heater On
05/31/2009 10:00:00 FILTER WHEEL The HIRS Filter Motor began stalling after 20009/151 10:00z. By the end of 2009/152, all HIRS data appears to be corrupted by the repeated stalling events.
05/10/2009 18:59:00   During the 18:59z Fairbanks pass on 5/10/2009 (Day 130) the Filter Wheel Period Monitor limits violated were Red High, Yellow Low on the Filter Wheel Sync, and Yellow High on the Filter Wheel Motor Current. After a few days of monitoring the HIRS began to quiet down ~05/14/2009
05/16/2005 09:00:00 FILTER WHEEL N15 HIRS filter wheel current abruptly dropped to approx 245 mAmps (nominal) vs the 390 that had been typical the past several weeks. Filter wheel synch and temperature also returned to nominal as did imagery. Subsystem and component upgraded to YELLOW status as a result.
04/18/2005 17:00:00 FILTER WHEEL No improvement in N15 HIRS after heater turnoff. Imagery still corrupted, filter wheel current still a or near record highs (390 mAmps), filter wheel synch always unlocked. Heater turnoff did not reduce record high filter wheel temp (32 degs C). Filter wheel and HIRS itself downgraded to RED status.
04/18/2005 14:25:00 FILTER WHEEL Turnoff HIRS filter heater to gauge filter motor response.
04/17/2005 13:02:47 FILTER WHEEL Major HIRS Filter motor anomaly 17 Apr 13:00Z. Filter wheel synch lost, period monitor erratic. Filter heater turned on 14:30Z.
12/11/2004 00:00:00 FILTER WHEEL Filter wheel current slowly returned back to and sustained nominal values (255 mAmps).
12/09/2004 12:00:00 FILTER WHEEL HIRS filter motor current surged over 40 mAmps to 280+ mAmps. Filter motor temp also rose and period monitor became more erratic. Filter wheel synch NOT lost at any time, no apparent imagery degradation. Began returning to normal (260)mA on10 Dec.
04/21/2004 13:45:00 FILTER WHEEL Turn off HIRS filter housing heater since filter motor current has improved.
04/19/2004 12:55:00 FILTER WHEEL Turn on HIRS filter housing heater because of filter motor current surge.
04/19/2004 02:00:00 FILTER WHEEL Major surge of filter wheel current and temps started at 02:00Z. Period monitor and filter wheel synch erratic. Imagery affected.
12/08/2003 18:25:00 FILTER WHEEL Turn off HIRS filter housing heater after filter wheel performance improved.
12/08/2003 15:00:00 FILTER WHEEL N15 HIRS filter wheel current improved significantly. Synch word and period monitor returned to (and sustained) nominal values.
12/06/2003 19:20:00 FILTER WHEEL Turn on HIRS filter housing heater in response to FW current surge.
12/06/2003 08:00:00 FILTER WHEEL HIRS filter motor current again surged as did filter motor temps. Period monitor and FW synch are again very erratic and imagery products are negatively impacted.
11/18/2003 16:15:00 FILTER WHEEL Turn off HIRS filter housing heater in response to HIRS improvement
11/16/2003 06:00:00 FILTER WHEEL N15 HIRS filter wheel current improved significantly. Synch word and period monitor returned to (and sustained) nominal values.
11/12/2003 15:15:00 FILTER WHEEL Turn on HIRS filter housing heater in response to filter motor current surge.
11/12/2003 12:00:00 FILTER WHEEL HIRS filter motor current experienced a significant runup as did filter motor temps. Period monitor and synch word were also very erratic.
07/29/2003 03:00:00 CHANNEL 17 HIRS Channels 13-19 showng increased space view noise with 14, 16 and 17 out of spec.
07/29/2003 03:00:00 CHANNEL 16 HIRS Channels 13-19 showng increased space view noise with 14, 16 and 17 out of spec.
07/29/2003 03:00:00 CHANNEL 14 HIRS Channels 13-19 showng increased space view noise with 14, 16 and 17 out of spec.
02/03/2003 05:00:00 FILTER WHEEL HIRS condition improved
02/02/2003 21:10:00 FILTER WHEEL Turn OFF HIRS filter housing heater
02/02/2003 00:35:00 FILTER WHEEL Turn on HIRS filter housing heater
02/01/2003 22:50:00 FILTER WHEEL HIRS had a substantial runup in filter motor current and temp. Period monitor and synch values were also anomalous and imagery was corrupted.
12/30/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
12/28/2002 00:10:00 FILTER WHEEL HIRS filter motor current surge
12/16/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
12/14/2002 00:00:00 FILTER WHEEL HIRS filter motor current surge
12/04/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
12/02/2002 07:00:00 FILTER WHEEL HIRS filter motor current surge
11/20/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
11/17/2002 18:00:00 FILTER WHEEL HIRS filter motor current surge
11/03/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
11/01/2002 00:00:00 FILTER WHEEL HIRS filter motor current surge
10/20/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
10/18/2002 00:00:00 FILTER WHEEL HIRS filter motor current surge
10/10/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
10/04/2002 12:00:00 FILTER WHEEL HIRS filter motor current surge
09/26/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
09/21/2002 00:00:00 FILTER WHEEL HIRS filter motor current surge
09/14/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
09/12/2002 17:00:00 FILTER WHEEL HIRS filter motor current surge
09/09/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
09/07/2002 00:01:00 FILTER WHEEL HIRS filter motor current surge
09/05/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
09/04/2002 02:56:00 FILTER WHEEL HIRS filter motor current surge
08/19/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
08/16/2002 22:39:00 FILTER WHEEL HIRS filter motor current surge
07/23/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
07/19/2002 15:00:00 FILTER WHEEL HIRS filter motor current surge
07/04/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
07/02/2002 01:00:00 FILTER WHEEL HIRS filter motor current surge
06/14/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
06/13/2002 00:01:00 FILTER WHEEL HIRS filter motor current surge
05/24/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
05/23/2002 00:00:00 FILTER WHEEL HIRS filter motor current surge
05/11/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
05/09/2002 21:00:00 FILTER WHEEL HIRS filter motor current surge
03/31/2002 23:00:00 FILTER WHEEL HIRS filter motor current runup and retreat, color changed to Yellow for inconsistency
03/24/2002 00:00:00 FILTER WHEEL HIRS filter motor current returned to nominal
03/23/2002 19:00:00 FILTER WHEEL HIRS filter motor current surge
02/26/2002 06:00:00 FILTER WHEEL HIRS Filter Motor Current runup, lasted through 27 Feb 01:00 then slowly returned to nominal on 1 Mar
02/13/2002 16:50:00 FILTER WHEEL HIRS Filter wheel heater turned OFF
02/12/2002 18:55:00 FILTER WHEEL HIRS Filter wheel heater turned ON
02/12/2002 13:45:00 FILTER WHEEL HIRS Filter motor current surged
01/31/2002 00:00:00 FILTER WHEEL Upgraded from Yellow because no problems observed since Oct 01
10/15/2001 04:30:00 FILTER WHEEL HIRS Filter motor current runup.
06/29/2001 00:00:00 FILTER WHEEL HIRS Filter motor heater turned off
06/27/2001 00:00:00 FILTER WHEEL HIRS Filter motor heater turned off because of erratic motor currents
02/07/2001 13:10:00 FILTER WHEEL HIRS filter wheel heater turned OFF. Channel 7 no longer being saturated.
01/13/2001 00:00:00 FILTER WHEEL HIRS miracle cure/self recovery
06/01/2000 00:00:00 FILTER WHEEL HIRS Filter wheel anomaly
12/28/1999 00:00:00 FILTER WHEEL HIRS Filter wheel anomaly
05/13/1998 16:24:00 CHANNEL 1 On Orbit
05/13/1998 16:24:00 CHANNEL 7 On Orbit
05/13/1998 16:24:00 CHANNEL 6 On Orbit
05/13/1998 16:24:00 CHANNEL 5 On Orbit
05/13/1998 16:24:00 CHANNEL 4 On Orbit
05/13/1998 16:24:00 CHANNEL 8 On Orbit
05/13/1998 16:24:00 CHANNEL 2 On Orbit
05/13/1998 16:24:00 CHANNEL 13 On Orbit
05/13/1998 16:24:00 ELECTRONICS On Orbit
05/13/1998 16:24:00 SCAN MOTOR On Orbit
05/13/1998 16:24:00 CHANNEL 3 On Orbit
05/13/1998 16:24:00 CHANNEL 9 On Orbit
05/13/1998 16:24:00 CHANNEL 10 On Orbit
05/13/1998 16:24:00 CHANNEL 12 On Orbit
05/13/1998 16:24:00 CHANNEL 14 On Orbit
05/13/1998 16:24:00 CHANNEL 15 On Orbit
05/13/1998 16:24:00 CHANNEL 16 On Orbit
05/13/1998 16:24:00 CHANNEL 17 On Orbit
05/13/1998 16:24:00 CHANNEL 18 On Orbit
05/13/1998 16:24:00 CHANNEL 19 On Orbit
05/13/1998 16:24:00 CHANNEL 20 On Orbit
05/13/1998 16:24:00 FILTER WHEEL On Orbit
05/13/1998 16:24:00 CHANNEL 11 On Orbit